# Pairwise Judgment Formulation for Semantic Embedding Model in Web Search

### v1

@@ 5 Hybrid Strategy
From Figures 2 and 3, a large performance discrepancy between any two atomic strategies is observed. Intuitively, combining any two of these atomic strategies will not bring better results than Clicked>Non-Examined since the low-quality training instances will ”contaminate” the result. However, through extensive empirical evaluation, we find that the intuition holds but with one exception, which results in the following hybrid strategy: Clicked>Non-Clicked, which is a combination of Clicked>Skipped and Clicked>Non-Examined.
Figure 4: Performance Comparison of Hybrid Strategies on Test-1 and Test-2
@@
Table 1: Pairwise Judgment Distribution of Atomic Strategies
@@
Based on the experimental results shown in Figure 4, we can see that Clicked>Non-Clicked slightly outperforms the best atomic strategy Clicked>Non-Examined on Test-1. In order to reveal the underlying reason, we present the distribution of the four atomic strategies. Based on the statistics shown in Table 1, we can see that no single strategy covers a majority of the pairs. Clicked>Skipped and Clicked>Non-Examined are responsible for 22.25% and 38.85% of the potential training data. Hence, relying on a single strategy such as Clicked>Non-Examined rules out many “fairly good” training instances and misses the chance of updating the embedding of many words. In the long run, the model using Clicked>Non-Clicked strategy sees more training instances achieves slightly better performance than the atomic strategies. The result shows that seeing more training instances is sometimes useful for training. Another possible explanation is that the testing data in Test-1 are essentially derived by Clicked>Non-Clicked strategy, hence, using Clicked>Non-Clicked has a better chance of getting a model with good fit for the testing data.
@@
In order to see whether the performance of Clicked>Non-Clicked is exaggerated by Test-1, we present the experimental result of Test-2 in Figure 4. We can see that the results of Test-2 is different from that of Test-1. In the beginning, Clicked>Non-Examined performs better than Clicked>Non-Clicked, indicating that high-quality pairwise judgments give a better start for training SEM. In the long run, the performance discrepancy of the two strategies diminishes. Clicked>Non-Examined and Clicked>Non-Clicked achieves similar performance when convergence is achieved.
Based on the observations in Section 4 and Section 5. The insights we obtained from such large-scale study are summarized as follows:
  1. 1.
@@
    The conventional LTR strategy for pairwise judgment formulation is not suitable for SEM. LTR aims to learn the weights of features in a ranking function. In contrast, the goal of SEM is to learn good representations of the features themselves.
  2. 2.
    The strategy Clicked>Non-Examined, which is rarely used in LTR, derives the best training data for SEM. Solely relying on this strategy significantly reduces the amount of training data for SEM and its performance is almost as good as using the strategy Clicked>Non-Clicked, which derives much more data for the training process to digest.
  3. 3.
@@
    When Clicked>Non-Examined is applied, adding pairwise judgments derived from Clicked>Skipped to form a hybrid heuristic causes no harm for training SEM, and may slightly improve its performance by extending the diversity of training instances.
6 Conclusions
@@
In this paper, we study the problem of formulating pairwise judgment for Semantic Embedding Model (SEM). We conduct an in-depth study of the information in the query log of a major search engine. Based on large-scale experiments, we quantitatively compare several strategies for pairwise judgment formulation. Experimental results verify the effectiveness of proposed strategies. More importantly, we reveal the fact that the pairwise judgment formulation for the neural network-based semantic embedding model is quite different from the pairwise LTR. Future work includes introducing more signals into the procedure of pairwise judgment formulation and investigating formulation strategies for SEM variants. The formulated pairwise judgment can also inspire downstream applications including topic discovery [9, 13, 16, 18], intent mining [10, 17], and user personalization [8, 12, 29].

### v2

@@ 6. Hybrid Strategy
In the previous section, the atomic strategies are investigated, and the Clicked¿Non-Examined is identified to provide the best performance in both tasks. A large performance discrepancy between any two atomic strategies is observed. Intuitively, combining these atomic strategies will not bring better results than Clicked¿Non-Examined since the low-quality training instances will “contaminate” the result. However, through extensive empirical evaluation, we find that the intuition holds but with one exception, which results in the following hybrid strategy: Clicked¿Non-Clicked, a combination of Clicked¿Skipped and Clicked¿Non-Examined.
Table 1. Distribution of Pairwise Judgment Dataset
@@
Clicked $>$ Non-Examined 38.87%
@@
Based on the experimental results shown in Figure 2(c), we can see that Clicked¿Non-Clicked slightly outperforms the best atomic strategy Clicked¿Non-Examined on Test-1. In order to reveal the underlying reason, we present the distribution of the four atomic strategies in Table 1 and the statistics suggest that no single strategy covers a majority of the pairs. Clicked¿Skipped and Clicked¿Non-Examined are responsible for 22.25% and 38.87% of the potential training data. Hence, relying on a single strategy such as Clicked¿Non-Examined rules out many “fairly good” training instances and misses the chance of updating the embedding of many words. In the long run, the model using Clicked¿Non-Clicked strategy sees more training instances, thus achieving slightly better performance than the atomic strategies. Another possible explanation is that the testing data in Test-1 are essentially derived by Clicked¿Non-Clicked strategy, hence resulting in a better chance of getting a model with a good fit for the testing data.
@@
In order to see whether the performance of Clicked¿Non-Clicked is exaggerated by Test-1, we also present the experimental result of Test-2 in Figure 2(c) and show that the results of Test-2 are different from those of Test-1. In the beginning, Clicked¿Non-Examined performs better than Clicked¿Non-Clicked, indicating that high-quality pairwise judgments give a better start for training SEM. In the long run, the performance discrepancy of the two strategies diminishes and results in similar performance when convergence is achieved.
Based on the observations from experimental results, the insights and best practices for pairwise judgment formulation are summarized as follows:
  1. (1)
@@
    The conventional LTR strategy for pairwise judgment formulation is not suitable for training SEM. LTR aims to learn the weights of features in a ranking function. In contrast, the goal of SEM is to learn good representations of the features themselves.
  2. (2)
@@
    When Clicked¿Non-Examined is applied, adding pairwise judgments derived from Clicked¿Skipped to form a hybrid heuristic may slightly improve its performance by extending the diversity of training instances.
  3. (3)
    The strategy Clicked¿Non-Examined, which is rarely used in LTR, derives the best training data for SEM. Solely relying on this strategy leads to a significantly smaller amount of training data (i.e., 38.87%) while maintaining the performance that is nearly as good as that achieved using the hybrid strategy Clicked¿Non-Clicked, which derives much more data (i.e., 61.10% as a combination of the two atomic strategies) for the training process.
7. Conclusions
@@
In this paper, we investigate the formulation of pairwise judgments for the Semantic Embedding Model (SEM) through an in-depth study of data from the query log of a major search engine. By conducting large-scale experiments, we quantitatively compare various strategies for pairwise judgment formulation and demonstrate the effectiveness of our proposed approaches. Notably, our findings reveal significant differences between the effective strategies for neural network-based SEM and traditional pairwise learning-to-rank (LTR) methods. Specifically, the strategy Clicked¿Non-Examined, although rarely utilized in LTR, proves to yield the highest quality training data for SEM. The hybrid strategy Clicked¿Non-Clicked, which encompasses a broader range of pairwise judgments, demonstrates slightly superior performance by providing a more extensive dataset for the model to learn from. Future work includes introducing more signals into the procedure of pairwise judgment formulation and investigating formulation strategies for SEM variants. The formulated pairwise judgment can also inspire downstream applications including topic discovery (Jiang et al., 2014; Jiang and Ng, 2013; Jiang et al., 2015b, 2023), intent mining (Jiang et al., 2016a; Jiang and Yang, 2016; Hong et al., 2024), and user personalization (Jiang et al., 2015a; Zhang et al., 2023).